OpenStudy (anonymous):

How do I simplify: (1/cosx)-(1/sinx)

OpenStudy (anonymous):

***corredction (1/sinx)/(1/cosx)

OpenStudy (ahsome):

Equation: \[\frac{1}{sinx}/\frac{1}{cosx}\] Right?

OpenStudy (ahsome):

@emilyhealy4?

OpenStudy (campbell_st):

when you divide by a fraction to rule is to flip the 2nd fraction and multiply so you get \[\frac{1}{\sin(x)} \times \frac{\cos(x)}{1} = \] and you'll find this is also a trig ratio

OpenStudy (ahsome):

Assuming it is: Turn the division sigh to a multiplication sign by flipping the numbers: \[\frac{1}{sinx}/\frac{1}{cosx}\]\[=\frac{1}{sinx}*\frac{cosx}{1}\] Simplify: \[\frac{cosx}{sinx}\] Cut the \(x\) \[\frac{cos}{sin}\]

OpenStudy (ahsome):

That is, if you want to cut the \(x\)